# MambaMorph: a Mamba-based Framework for Medical MR-CT Deformable Registration

### v2

@@ 4.1.2 Model selection and implementation
To validate our MambaMorph’s superiority in multi-modal deformable registration, we compare it with two representative DL-based methods (i.e. VoxelMorph and TransMorph). In quantitative comparison (Tab. 1), we evaluate the complete MambaMorph and the original MambaMorph (MambaMorph). Specifically, MambaMorph is equipped with Mamba-based registration module, feature extractor and integration layer. For fair comparison, MambaMorph is almost identical to TransMorph, which only replaces Swin Transformer block with Mamba block. In ablation study (Tab. 2), we prove the advantage of introducing fine-grained but simple feature extractor to the three registration modules. The feature extractor is a 2-layer depth UNet whose channel numbers are 16 at all layers. Notably, the MambaMorph without feature extractor includes integration layer and the MambaMorph with feature extractor is the same to the complete MambaMorph at Tab. 1.
@@ 4.2.1 Quantitative comparison
From Tab. 1, we can see that our MambaMorph surpasses TransMorph in nearly all aspects except for diffeomorphism. Only with a single substitution, can our registration module be more accurate, more lightweight and faster than TransMorph. These experiments demonstrate the excellence of Mamba in contrast to Transformer, which may attribute to Mamba’s outstanding long-range modeling ability and its nearly linear complexity. Thanks to the efficient Mamba block, our MambaMorph is about 8% higher on Dice and 0.7 lower on HD95 than TransMorph, consuming more inference time and memory slightly. Due to superior registration performance, acceptable inference time and low memory consumption, we assume that Mamba block is more suitable than Transformer in practical deformable registration.

### v5

@@ 4.1.2 Model selection and implementation
To validate our MambaMorph’s superiority in multi-modal deformable registration, we compare it with two representative DL-based methods (i.e. VoxelMorph and TransMorph). In quantitative comparison (Tab. 1), we evaluate the complete MambaMorph and the original MambaMorph (MambaMorphori). Specifically, MambaMorph is equipped with Mamba-based registration module, feature extractor and integration layer. For fair comparison, MambaMorphori is almost identical to TransMorph, which only replaces Swin Transformer block with Mamba block. In ablation study (Tab. 2), we prove the advantage of introducing fine-grained but simple feature extractor to the three registration modules. The feature extractor is a 2-layer depth UNet whose channel numbers are 16 at all layers. Notably, the MambaMorph without feature extractor includes integration layer and the MambaMorph with feature extractor is the same to the complete MambaMorph at Tab. 1.
@@ 4.2.1 Quantitative comparison
From Tab. 1, we can see that our MambaMorphori surpasses TransMorph in nearly all aspects except for diffeomorphism. Only with a single substitution, can our registration module be more accurate, more lightweight and faster than TransMorph. These experiments demonstrate the excellence of Mamba in contrast to Transformer, which may attribute to Mamba’s outstanding long-range modeling ability and its nearly linear complexity. Thanks to the efficient Mamba block, our MambaMorph is about 8% higher on Dice and 0.7 lower on HD95 than TransMorph, consuming more inference time and memory slightly. Due to superior registration performance, acceptable inference time and low memory consumption, we assume that Mamba block is more suitable than Transformer in practical deformable registration.